Find the value of z so that the area under the standard normal curve to the right of z is 0.3317. Round to four decimal places.
Answer #1

Here we want to find z such that P( Z > z ) = 0.3317

Let P( Z < z)= 1 - P(Z > z ) = 1 - 0.3317 = 0.6683

Let's use Excel:

z = "=NORMSINV(0.6683)" = 0.4352

So the value of z corresponding to the probability 0.6683 is 0.4352